# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 2.1 What is Robustness in LLMs
  • Stability under Disruption: Robustness is fundamentally about the stability of LLM performance and behaviour when faced with various forms of disruption, including unseen scenarios, different types of attacks, and noise in the input (Guo et al., 2023). It implies predictability even when conditions deviate from the ideal.
  • Performance Maintenance under Distribution Shifts (Du et al., 2022): Robustness entails maintaining performance levels when the distribution of deployment data differs from the training data (OOD generalization). This is crucial as real-world data rarely perfectly matches training sets.
  • Resilience to Input Variations: A robust LLM should maintain consistent outputs despite variations in the input that preserve the core meaning or intent. This includes resilience to natural language variations like morphological changes, typos, paraphrasing, and syntactic restructuring (Singh et al., 2024; Mahaut et al., 2024; Sarker et al., 2024). It also encompasses stability when confronted with unforeseen or malformed prompts (Ailem et al., 2024).
  • Consistency and Reliability of Outputs: Robustness requires the model to consistently generate outputs that are accurate, reliable, and unbiased across diverse scenarios (Lad et al., 2024; Aljohani et al., 2025). This involves minimizing errors, factual inaccuracies (hallucinations), and harmful biases, even under challenging conditions.
  • Adherence to Intended Behaviour: A robust model should adhere to its intended function and safety constraints. This includes following legitimate instructions correctly while resisting malicious manipulations like prompt injection attacks (Li et al., 2024d) or jailbreaks designed to elicit harmful content.
@@ 2.2 Key Dimensions of Robustness
LLM robustness is not a monolithic property but rather a collection of related capabilities. To better understand and address the challenge, it is useful to delineate the key facets of LLM robustness, recognizing that these categories often overlap and interact. Figure 2 provides an overview of this overlapping nature, illustrating how progress in one dimension can amplify or undermine robustness in others due to their inherent interdependencies
  1. 1.
    Resilience to Adversarial Attacks (Input Perturbations): This dimension concerns the model’s ability to withstand inputs that have been intentionally manipulated by an adversary to cause specific failures (Chacko et al., 2024). These perturbations are often small and designed to be imperceptible or semantically plausible to humans but exploit model vulnerabilities.
    • Appearance: Misclassification in tasks like sentiment analysis (Yang et al., 2024d), generation of harmful, toxic, or biased content, bypassing safety alignments (jailbreaking) (Tao et al., 2024), or general performance degradation.
@@
    • Attack Types: Attacks on LLMs can target varying levels of input granularity: character-level perturbations (e.g., HotFlip), word-level modifications (e.g., synonym swapping), or sentence-level alterations (e.g., inserting distracting phrases) (Yang et al., 2024d). Beyond these, typographical errors have been used as adversarial tools as in Gan et al. (2024). Attacks are further categorized by threat model - white-box attacks assume access to model internals (e.g., parameters or logits), whereas black-box attacks interact solely via inputs and outputs. Prominent techniques include the Greedy Coordinate Gradient (GCG), AutoDAN, and PAIR (Xhonneux et al., 2024), with some methods focusing on perturbing continuous embeddings rather than discrete tokens. Multimodal LLMs (MLLMs) also face compounded risks, as adversaries exploit vulnerabilities both within individual modalities (e.g., image or text) and across modality interactions (Jiang et al., 2025).
    • Cause: Exploitation of model vulnerabilities, including reliance on spurious correlations and brittle statistical dependencies.
    Figure 2: Illustration of challenges in achieving comprehensive LLM robustness due to interconnected and potentially conflicting dimensions.
  2. 2.
    Out-of-Distribution (OOD) Generalization: This refers to the model’s ability to maintain performance when evaluated on data drawn from a distribution different from the one used for training. Real-world data distributions often shift over time or across contexts, making OOD generalization crucial for practical deployment.
    • Appearance: Significant performance drop occurs in LLMs when evaluated on out-of-distribution (OOD) test sets compared to in-distribution (ID) benchmarks (Du et al., 2022). This degradation arises from two primary failure modes: covariate shift (divergence in input features) or concept shift (changes in input-output relationships) (Yuan et al., 2023; Nagarajan et al., 2024). Specific OOD challenges include handling unfamiliar knowledge domains (OOD Knowledge), stylistic variations (OOD Style), or multilingual/code-mixed inputs (OOD Language).
    • Cause: Model’s tendency to exploit spurious correlations or training data biases rather than learning generalizable invariant features.
  3. 3.
    Sensitivity to Prompt Variations: This dimension measures how stable the model’s output is when the input prompt or instruction is rephrased or slightly altered while preserving the original semantic meaning (Yang et al., 2024d). Humans are generally resilient to such variations, but LLMs can be surprisingly sensitive (Agrawal et al., 2025).
    • Appearance: Drastic changes in output quality, correctness, or consistency even with minor prompt modifications like paraphrasing, synonym replacement, word insertion/deletion, or reordering of elements (e.g., multiple-choice options).
    • Cause: Overfitting to specific instruction formats seen during training or reliance on superficial cues in the prompt.
  4. 4.
    Handling Noisy or Corrupted Input: This concerns the model’s resilience to unforeseen mistakes or flaws commonly found in real-world user inputs.
    • Appearance: LLMs often struggle to process inputs containing textual noise such as typos, misspellings, grammatical errors, or irregular punctuation. This includes errors introduced by automated processes (e.g., voice-to-text transcription flaws or OCR mistakes). Sensitivity extends to noise within training or fine-tuning data itself, which can degrade robustness by reinforcing brittle input-output mappings (Luo et al., 2024).
    • Cause: Reinforcement of brittle input-output mappings from noisy training/fine-tuning data or the context-aware nature of LLMs’ design, making them susceptible to distraction by irrelevant information.
  5. 5.
    Consistency and Reliability of Outputs: This focuses on the dependability of the model’s generations, ensuring outputs are stable for similar inputs and free from self-contradiction or factual inaccuracies (hallucinations) (Zhou et al., 2024b).
    • Appearance: Generating different answers to semantically identical questions posed slightly differently (Yang et al., 2025b). Further, producing factually incorrect or nonsensical information (hallucination), which can be exacerbated by robustness issues or outputting contradictory statements within the same response or across closely related prompts.
    • Cause: Lack of stable internal representation or deterministic reasoning or over-reliance on statistical fluency at the expense of global factual accuracy or logical soundness.
  6. 6.
    Fairness under Stress: The ability to maintain fairness and avoid amplifying societal biases (e.g., related to gender, race, or other protected attributes) when presented with adversarial prompts or contexts specifically designed to induce biased behaviour (Jung et al., 2025).
    • Appearance: Performance disparities across diverse socio-demographic and geographic contexts. For instance, LLMs have been observed to consistently outperform on U.S. datasets compared to Chilean datasets, indicating a geographical bias in their performance. Furthermore, the influence of protected attributes on prediction accuracy varies by region in the U.S., political identity and race significantly affect accuracy, whereas in Chile, gender, education, and religious affiliation play more pronounced roles.
    • Cause: Flaws in training data leading to representational imbalance and skewed learned distributions.
  7. 7.
    Task-Specific Robustness: Robustness considerations tailored to the demands of specific downstream tasks, such as maintaining logical coherence in reasoning tasks or generating functionally correct and secure code (Bao et al., 2025; Li et al., 2025d).
    • Appearance: Significant challenges in logical reasoning, failure in complex logical problems (e.g., low accuracy on FOLIO (Han et al., 2024)), violation of logical consistency. Furthermore, the difficulties highlighted in solver-aided approaches, such as information loss, translation errors, and limited translation capacities, lead to a significant “symbolic gap”.
    • Cause: Inherent limitations in formal reasoning and symbolic manipulation, and challenges in robustly bridging natural language and formal logic.
These dimensions are often interconnected. For example, a model susceptible to spurious correlations (an OOD issue) might also be more vulnerable to adversarial attacks that exploit those correlations, or more sensitive to prompt variations that alter the presence of those spurious cues. Further, inputs that are OOD may push a model beyond its knowledge boundaries, thereby increasing the likelihood of hallucination. Similarly, improving robustness against noise might also enhance consistency. However, it is also conceivable that improving one dimension might negatively impact another. For instance, alignment techniques aimed at improving fairness (a form of robustness) might inadvertently reduce OOD generalization (the “alignment tax”) (Lin et al., 2024b). Figure 3 highlights these dependencies and trade-offs, emphasizing the need for balanced strategies to advance robustness across interdependent dimensions.
@@
Figure 3: A conceptual visualization of the interdependent dimensions critical to LLM robustness, highlighting the synergies and tensions between improving OOD generalization, noise resilience, output consistency, and fairness. For instance, enhancing consistency (e.g., reducing sensitivity to prompt variations) lowers vulnerability to adversarial attacks, while improving OOD generalization reduces hallucination risks and may indirectly reinforce consistency. However, over-optimizing one dimension (e.g., rigidly enforcing consistency) could constrain adaptability to novel inputs. This visualization demonstrates how these factors interconnect, highlighting the challenge of achieving well-balanced, comprehensive robustness across all dimensions.
3 Sources of Robustness
@@
Understanding why LLMs lack robustness requires examining the contributing factors across different stages of their lifecycle, from data collection to training and deployment. These issues can be broadly categorized as stemming from the data, the model itself, or the learning process. Table 1 provides an overview of these sources. These failures in robustness can originate mainly from:
@@
  • Inference-Related:
@@ 3.1.1 Spurious Correlations
This is arguably one of the most significant causes of robustness failures, particularly concerning OOD generalization (Du et al., 2022). Instead of learning the intended, often complex, causal relationships or concepts, models learn to rely on simpler, superficial patterns (shortcuts) that happen to correlate with the correct output in the training data but do not hold universally. When these spurious correlations are absent or altered in new inputs (e.g., OOD data, perturbed examples), the model’s performance collapses. Examples include:
  • Lexical Bias: Relying on specific words (e.g., negation words like ‘no’, ‘never’ strongly predicting ‘contradiction’ in Natural Language Inference (NLI) tasks).
  • Overlap Bias: Exploiting word overlap between premise and hypothesis in NLI.
  • Positional Bias: Predicting answers based on their typical position in the training context (e.g., always in the k-th sentence for QA).
  • Style Bias: Associating certain writing styles with specific labels.
  • Heuristics in Reasoning: Using memorised rules-of-thumb or simple numerical patterns instead of robust algorithms (Li et al., 2024b) for tasks like arithmetic.
3.1.2 Dataset Biases and Anomalies
LLMs often inherit shortcomings from systematic biases and unintended statistical cues in their training data. Training data inevitably reflects societal biases present in the source text (often large web-scraped corpora). LLMs can learn and amplify these biases related to gender, race, culture, religion, etc., leading to unfair, discriminatory, or non-robust outputs when prompted about sensitive topics. Evaluating fairness under prompts designed to induce bias reveals these vulnerabilities (Jung et al., 2025). For instance, crowd-sourced datasets, used widely for fine-tuning and evaluation, frequently contain superficial annotation patterns, where human annotators unintentionally introduce simplistic labeling strategies that models exploit as shortcuts. These patterns create misleading statistical associations (e.g., keyword matching) rather than reflecting genuine task understanding. Even benchmark design can amplify biases, as models may overfit to variations in test data construction rather than solving tasks as intended (Ailem et al., 2024). Together, these data biases and methodological flaws perpetuate robustness failures across training, evaluation, and deployment.
3.1.3 Data Poisoning/Backdoors
Adversaries can intentionally inject malicious examples into the training data (or fine-tuning data) to create hidden backdoors. These backdoors can be triggered by specific inputs at inference time to cause targeted misbehaviour or compromise model safety, often without degrading general performance noticeably, specifically in fine-tuned models (Aljohani et al., 2025).
3.1.4 Sensitivity to Input Variations
  • Prompt Sensitivity: Minor changes in wording, punctuation, or formatting of prompts can lead to significantly different and sometimes incorrect outputs. Models might solve a task correctly with one phrasing but fail with equivalent rewordings (Mahaut et al., 2024). Even minor typos can degrade the performance (Liu et al., 2024b).
  • Instruction Sensitivity: Performance can vary drastically based on how instructions are phrased or structured within the prompt. Counterintuitively, models specifically fine-tuned for a domain might become more sensitive to instruction variations than general-purpose models (Yan et al., 2024a; Aljohani et al., 2025).
  • Noise Sensitivity: Models struggle with naturally occurring noise (spelling/grammar errors, OCR/ASR anomalies) (Singh et al., 2024) and can be easily distracted by irrelevant information or deliberately injected noise (Zhou et al., 2024d).

### v2

@@ 2.1 What is Robustness in LLMs
  • Stability under Disruption: Robustness is fundamentally about the stability of LLM performance and behaviour when faced with various forms of disruption, including unseen scenarios, different types of attacks, and noise in the input (guo2023evaluatinglargelanguagemodels). It implies predictability even when conditions deviate from the ideal.
  • Consistency and Reliability of Outputs: Robustness requires the model to consistently generate outputs that are accurate, reliable, and unbiased across diverse scenarios (lad2024remarkablerobustnessllmsstages; aljohani2025comprehensivesurveytrustworthinesslarge). This involves minimizing errors, factual inaccuracies (hallucinations), and harmful biases, even under challenging conditions.
  • Adherence to Intended Behaviour: A robust model should adhere to its intended function and safety constraints. This includes following legitimate instructions correctly while resisting malicious manipulations, such as prompt injection attacks (li-etal-2024-evaluating-instruction), (which involves performing unintended actions by combining adversarially designed malicious input with the model’s instructions) or backdoor attacks (cai2022badprompt; chen2021badnl).
  • Resilience to Input Variations: A robust LLM should maintain consistent outputs despite variations in the input that preserve the core meaning or intent. This includes resilience to natural language variations like morphological changes, typos, paraphrasing, and syntactic restructuring (singh2024robustnessllmsperturbationstext; Mahaut_2024; sarker2024syntacticrobustnessllmbasedcode). It also encompasses stability when confronted with unforeseen or malformed prompts (ailem2024examiningrobustnessllmevaluation).
  • Performance Maintenance under Distribution Shifts (du2022robustness): Robustness entails maintaining performance levels when the distribution of deployment data differs from the training data (OOD generalization). This is crucial as real-world data rarely perfectly matches training sets.
@@ 2.2 Key Dimensions of Robustness
LLM robustness is not a monolithic property but rather a collection of related capabilities. To better understand and address the challenge, it is useful to delineate the key facets of LLM robustness, recognizing that these categories often overlap and interact. Figure 3 provides an overview of this overlapping nature, illustrating how progress in one dimension can amplify or undermine robustness in others due to their inherent interdependencies.
2.2.1 Resilience to Adversarial Attacks
It refers to a model’s ability to withstand inputs that have been intentionally manipulated by an adversary to trigger specific failures (chacko2024adversarialattackslargelanguage). Such perturbations are typically small, often imperceptible or semantically plausible to humans, yet they exploit latent vulnerabilities in the model. These attacks commonly manifest as misclassifications in classification tasks (yang2024assessingadversarialrobustnesslarge), the generation of harmful or toxic content, or the circumvention of safety alignments through jailbreaks (a class of prompt injections designed to override safety filters and moderation policies) (tao2024robustnesslargelanguagemodels). In broader cases, adversarial perturbations can simply degrade model performance without obvious failure.
@@ 2.2.1 Resilience to Adversarial Attacks
Attacks on LLMs vary in their granularity. At the input level, adversaries can manipulate characters (e.g., HotFlip), words (e.g., synonym substitution), or entire sentences (e.g., insertion of distracting phrases) (yang2024assessingadversarialrobustnesslarge). More subtle perturbations such as typographical errors have also been shown to degrade reasoning robustness (gan2024reasoningrobustnessllmsadversarial). Attacks further differ depending on the threat model: white-box attacks assume access to internal parameters or gradients, while black-box attacks rely solely on interacting with inputs and outputs. Prominent techniques include Greedy Coordinate Gradient (GCG), AutoDAN, and PAIR (xhonneux2024efficientadversarialtrainingllms), with some methods targeting continuous embeddings rather than discrete tokens. In the case of multimodal LLMs (MLLMs), adversaries can exploit vulnerabilities within individual modalities (e.g., image or text) or across modality interactions, compounding the risk (jiang2025surveyadversarialrobustnessmultimodal).
At their core, adversarial attacks succeed by exploiting brittle statistical dependencies and spurious correlations learned during training, exposing the fragility of LLMs to small but strategically chosen perturbations.
2.2.2 OOD Genralization
It concerns the ability of a model to sustain performance when faced with inputs drawn from distributions different from those encountered during training. Real-world deployment frequently involves such distribution shifts, arising from temporal drift, novel domains, or different user populations. When tested on OOD datasets, LLMs often experience significant performance degradation relative to in-distribution (ID) benchmarks (du2022robustness). Failures commonly emerge from covariate shift, where input features differ, or concept shift, where input–output mappings change (yuan2023revisiting; nagarajan2024understandingfailuremodesoutofdistribution).
These challenges surface in multiple forms: difficulties with unfamiliar knowledge domains, stylistic shifts in text, or multilingual and code-mixed settings. Such failures reflect a tendency to rely on spurious correlations and biases present in training data, rather than learning generalizable, invariant features.
2.2.3 Sensitivity to Prompt Variations
It captures how much a model’s output changes when the input instruction is rephrased or slightly altered while retaining its semantic intent (yang2024assessingadversarialrobustnesslarge). Unlike humans, who are generally resilient to paraphrasing or minor structural shifts, LLMs can exhibit drastic changes in correctness, coherence, or consistency under such conditions (agrawal2025enhancingllmrobustnessperturbed).
Even simple alterations such as synonym replacement, insertion or deletion of words, or reordering of answer options can destabilize performance. This fragility often stems from overfitting to prompt formats encountered during training and reliance on superficial cues rather than deeper semantic understanding. Reducing sensitivity to prompt variation is therefore essential for achieving robustness in instruction-following scenarios.
2.2.4 Handling Noisy or Corrected Input
It refers to a model’s capacity to process data containing imperfections commonly found in real-world use. Such noise may include typographical errors, misspellings, irregular grammar, or extraneous punctuation. Automated pipelines introduce additional errors, e.g., transcription mistakes from voice-to-text systems or inaccuracies from OCR (Optical Character Recognition) processes.
Models trained or fine-tuned on noisy data are further at risk, as they may learn brittle mappings that amplify fragility rather than tolerance (luo2024robustftrobustsupervisedfinetuning). In practice, noisy or corrupted inputs can distract the model and degrade downstream task performance, underscoring the need for architectures and training regimes that explicitly account for the messiness of real-world data.
2.2.5 Fairness Under Stress
It highlights a model’s ability to uphold fairness and avoid amplifying societal biases when confronted with adversarial prompts or challenging contexts (jung2025flexbenchmarkevaluatingrobustness). Failures in this dimension manifest as disparities in prediction quality across socio-demographic or geographic groups. For instance, LLMs often perform better on U.S.-based datasets than on Chilean ones, reflecting geographical bias. In the U.S., attributes such as race and political identity influence accuracy, whereas in Chile, gender, education, and religious affiliation play stronger roles (abeliuk2025fairnessllmgeneratedsurveys; Qu_Wang_2024).
These disparities typically originate from imbalances in training data, where certain populations are underrepresented or misrepresented, resulting in skewed learned distributions. Addressing this requires approaches that ensure equitable performance across diverse populations, even under adversarial or distributionally shifted conditions.
2.2.6 Consistency and Reliability of Outputs
Consistency and reliability pertain to the stability and trustworthiness of a model’s generations across semantically similar inputs. Failures of consistency include producing contradictory answers to paraphrased queries or generating factually incorrect, nonsensical, or hallucinated content (zhou2024trustworthinessretrievalaugmentedgenerationsystems; yang2025enhancingsemanticconsistencylarge).
These issues arise from a lack of stable internal representations and an overemphasis on local statistical fluency at the expense of factual soundness or logical coherence. As such, inconsistency is closely tied to other robustness dimensions, including OOD generalization and adversarial susceptibility. Ensuring consistency is not only a matter of reliability but also a prerequisite for deploying LLMs in domains where factual accuracy is critical.
@@ 2.2.6 Consistency and Reliability of Outputs
Figure 3: A conceptual visualization of the interdependent dimensions critical to LLM robustness, highlighting the synergies and tensions between improving OOD generalization, noise resilience, output consistency, and fairness, and so highlighting the challenge of achieving well-balanced, comprehensive robustness across all dimensions.
2.2.7 Task-Specific Robustness
Task-specific robustness addresses the challenges LLMs face in meeting the unique demands of particular downstream applications. In reasoning-intensive tasks, models often struggle to maintain logical consistency, as seen in low performance on benchmarks such as FOLIO (han2024folionaturallanguagereasoning). In coding applications, failures include the generation of functionally incorrect or insecure code (bao2025assessingenhancingrobustnesslarge).
Such difficulties stem from the symbolic gap between natural language and formal logic or programming languages, as well as from translation errors and capacity limits in solver-aided methods. These weaknesses highlight that robustness cannot be fully captured by general benchmarks alone but must be evaluated in domain-specific contexts where performance reliability is crucial.
These dimensions are often interconnected. A model prone to spurious correlations (an OOD issue) is often more vulnerable to adversarial attacks that exploit those patterns and more sensitive to prompt variations that shift these cues. Enhancing consistency (e.g., reducing prompt sensitivity) can lower adversarial vulnerability, while stronger OOD generalization reduces hallucination risks and may indirectly reinforce consistency. Robustness to noise may also strengthen consistency. At the same time, progress in one dimension can create trade-offs in another. For example, enforcing consistency too rigidly may limit adaptability to novel inputs, and alignment methods aimed at improving fairness may weaken OOD generalization (also known as the “alignment tax”) (lin2024mitigatingalignmenttaxrlhf). Moreover, OOD inputs can push models beyond their knowledge boundaries, increasing hallucination risks. Overall, robustness requires balanced strategies, as over-optimizing a single dimension can undermine others. Figure 3 illustrates these dependencies and trade-offs across interdependent dimensions.
3 Sources of Non-Robustness
@@
Understanding why LLMs lack robustness requires examining the contributing factors across different stages of their lifecycle, from data collection to training and deployment. Figure 4 illustrates some of the failure cases by giving some real-time examples. Further, these issues can be broadly categorized as stemming from the data, the model itself, or the learning process. Table 2 provides an overview of these sources. These failures in robustness can originate mainly from:
@@
  • Inference-Related: Vulnerabilities that arise during the model’s deployment or use, such as issues from the choice of decoding strategies, reliance on external retrieval systems, or exposure to distribution shifts and adversarial inputs at inference time.
@@ 3.1.1 Spurious Correlations
This is arguably one of the most significant causes of robustness failures, particularly concerning OOD generalization (du2022robustness; mcmilin2022selection). Instead of learning the intended, often complex, causal relationships or concepts, models learn to rely on simpler, superficial patterns (shortcuts) that happen to correlate with the correct output in the training data but do not hold universally. When these spurious correlations are absent or altered in new inputs (e.g., OOD data, perturbed examples), the model’s performance collapses. Examples include:
  • Lexical Bias: Relying on specific words (e.g., negation words like ‘no’, ‘never’ strongly predicting ‘contradiction’ in Natural Language Inference (NLI) tasks) BiCapsHATE; du2023shortcutlearninglargelanguage.
  • Overlap Bias: Exploiting word overlap between premise and hypothesis in NLI li2025overlap; chandna2025dissectingbiasllmsmechanistic.
  • Positional Bias: Predicting answers based on their typical position in the training context (e.g., always in the k-th sentence for QA) positionwang2023largelanguagemodelsfair; positionshi2025judgingjudgessystematicstudy.
  • Style Bias: Associating certain writing styles with specific labels Cao_2025style; malik-etal-2024-empirical.
  • Heuristics in Reasoning: Using memorised rules-of-thumb or simple numerical patterns instead of robust algorithms (li-etal-2024-gsm) for tasks like arithmetic.
@@
3.1.2 Dataset Biases and Anomalies
LLMs often inherit shortcomings from systematic biases and unintended statistical cues in their training data. Training data inevitably reflects societal biases present in the source text (often large web-scraped corpora). LLMs can learn and amplify these biases related to gender, race, culture, religion, etc., leading to unfair, discriminatory, or non-robust outputs when prompted about sensitive topics. Evaluating fairness under prompts designed to induce bias reveals these vulnerabilities (jung2025flexbenchmarkevaluatingrobustness). For instance, crowd-sourced datasets, used widely for fine-tuning and evaluation, frequently contain superficial annotation patterns, where human annotators unintentionally introduce simplistic labeling strategies that models exploit as shortcuts. These patterns create misleading statistical associations (e.g., keyword matching) rather than reflecting genuine task understanding. Even benchmark design can amplify biases, as models may overfit to variations in test data construction rather than solving tasks as intended (ailem2024examiningrobustnessllmevaluation). Together, these data biases and methodological flaws perpetuate robustness failures across training, evaluation, and deployment.
3.1.3 Data Poisoning/Backdoors
Adversaries can intentionally inject malicious examples into the training data (or fine-tuning data) to create hidden backdoors. These backdoors can be triggered by specific inputs at inference time to cause targeted misbehaviour or compromise model safety, often without degrading general performance noticeably, specifically in fine-tuned models (aljohani2025comprehensivesurveytrustworthinesslarge).
3.1.4 Sensitivity to Input Variations
  • Prompt Sensitivity: Minor changes in wording, punctuation, or formatting of prompts can lead to significantly different and sometimes incorrect outputs. Models might solve a task correctly with one phrasing but fail with equivalent rewordings (Mahaut_2024). Even minor typos can degrade the performance (liu2024trustworthyllmssurveyguideline).
  • Instruction Sensitivity: Performance can vary drastically based on how instructions are phrased or structured within the prompt. Counterintuitively, models specifically fine-tuned for a domain might become more sensitive to instruction variations than general-purpose models (yan2024contrastiveinstructiontuning; aljohani2025comprehensivesurveytrustworthinesslarge).
  • Noise Sensitivity: Models struggle with naturally occurring noise (spelling/grammar errors, OCR/ASR anomalies) (singh2024robustnessllmsperturbationstext) and can be easily distracted by irrelevant information or deliberately injected noise (zhou2024languagemodelsperformrobust).